CH 3 Quiz ACG 2071 Spring 2019 Quiz: CH 3 Quiz This Question: 1 pt Joseph Pellicano & Time Remaining: 00:43.06 Submit Quiz 4 of 20 (4 complete) This Quiz: 20 pts possible ot Showboat Corporation had actual manufacturing overhead costs for the most recent year of $29,900. Manufacturing overhead is allocated using a predetermined manufacturing overhead rate of $4.25 per direct labor hour. Direct labor cost is $28 per hour. At the end of the year, Cabaret Corporation found it had overallocated manufacturing overhead by $1,650. How many direct labor hours were worked in total during the year? (Round your final answer to the nearest dollar.) O A 7.035 OB. 6,647 O C. 7.424 O D. 7.013
Answer #1
Over-allocated manufacturing overhead = Budgeted manufacturing overhead - Actual manufacturing overhead
1650 = Budgeted manufacturing overhead - 29900
Budgeted manufacturing overhead = 29900+1650 = 31550
Manufacturing overhead rate = Budgeted manufacturing overhead costs / Budgeted labour hours
4.25 = 31550 / Budgeted labour hours
Budgeted labour hours = 31550/4.25 = 7424 hours